# Maintaining Distributed Data Structures in Dynamic Peer-to-Peer Networks

### v1

@@ 2 Solution Architecture
Let us briefly recall skip lists [68], that are randomized data structures organized as a tower of increasingly sparse linked lists. Level 00 of a skip list is a classical linked list of all nodes in increasing order by key/ID. For each $i$ such that $i>0$, each node in level $i-1$ appears in level $i$ independently with some fixed probability $p$. The top lists act as “express lanes” that allow the sequence of nodes to be quickly traversed. Searching for a node with a particular key involves searching first in the highest level, and repeatedly dropping down a level whenever it becomes clear that the node is not in the current one. By backtracking on the search path it is possible to shows that no more than $\frac{1}{1-p}$ nodes are searched on average per level, giving an average search time of $\mathcal{O}(\log n)$ (see Appendix B for useful properties of randomized skip lists). We refer to the height of a skip list $\mathcal{L}$ (see Figure 2) as the maximum $h$ such that Level $h$ is not empty. Given a node $v$ in the skip list $\mathcal{L}$ with $n_{\mathcal{L}}$ elements, we use $N_{\mathcal{L}}^{(R,\ell)}(v)$ and $N_{\mathcal{L}}^{(L,\ell)}(v)$ to indicate $v$’s right and left neighbors at level $\ell$ respectively. Moreover, when the direction and the levels are not specified we refer to the overall set of neighbors of a node in a skip list, formally we refer to $N_{\mathcal{L}}(v)=\bigcup_{\ell\geq 0}(N_{\mathcal{L}}^{(L,\ell)}(v)\cup N_{\mathcal{L}}^{(R,\ell)}(v))$. Furthermore, we define $\ell_{\max}^{\mathcal{L}}(v)$ for a node $v\in\mathcal{L}$ to be its maximum height in the skip list [^5] We omit the superscript when the skip list is clear from the context. and $\ell_{\max}(\mathcal{L})=\max_{v\in\mathcal{L}}\ell_{\max}^{\mathcal{L}}(v)$ to be the overall height of the skip list.
Figure 2: Skip list with $n=6$ nodes and $3$ levels.
@@
Moreover, in all skip lists, when a node exits the system, it is operated by a selected group of nodes. In the course of our algorithm description, if a node $u$ is required to perform some operation, but is no longer in the system, then its replacement node(s) will perform that operation on its behalf. Note further that some of the replacement nodes themselves may need to be replaced. Such replacement nodes will continue to represent $u$. The protocol assumes a short ($\Theta(\log n)$ round) initial “bootstrap” phase, where there is no churn [^6] Without a bootstrap phase, it is easy to show that the adversary can partition the network into large pieces, with no chance of forming even a connected graph. and it initializes the underlying network. More precisely, the bootstrap is divided in two sub-phases in which we (i) build the underlying churn resilient network described in Section 2.1 in $\mathcal{O}(\log n)$ rounds and, (ii) we build the skip lists data structures $\mathcal{L}$ and $\mathcal{C}$ (initially $\mathcal{L}=\mathcal{C}$) using the $\mathcal{O}(\log n)$ rounds technique described in Section 2.4; after this, the adversary is free to exercise its power to add or delete nodes up to the churn limit and the network will undergo a continuous maintenance process. The overall maintenance of the dynamic distributed data structure goes through cycles. Each cycle $c\geq 1$ in Algorithm 1 is comprised of four phases. Without loss of generality assume that initially $\mathcal{L}$ and $\mathcal{C}$ are the same. We use the notation $\mathcal{L}^{(c)}$ (resp. $\mathcal{S}^{(c)},\mathcal{C}^{(c)},\mathcal{B}^{(c)}$) to indicate the network $\mathcal{L}$ (resp. $\mathcal{S},\mathcal{C},\mathcal{B}$) during the cycle $c\in\mathbb{N}$, when it is clear from the context we omit the superscript to maintain a cleaner exposition.
@@
7      Phase 3 (Merge): The buffer $\mathcal{B}^{(c)}$ created in Phase $2$ is merged into $\mathcal{C}^{(c)}$, i.e., $\mathcal{C}^{(c+1)}\leftarrow\mathcal{C}^{(c)}\cup\mathcal{B}^{(c)}$.
@@ 2.1 The Spartan network
A useful technique to build and maintain a stable overlay network that is resilient to a high amount of adversarial churn is to construct and maintain a network of (small-sized) committees [7, 9, 29, 13]. A committee is a clique of small ($\Theta(\log n)$) size composed of essentially “random” nodes. A committee can be efficiently constructed, and more importantly, maintained under large churn. Moreover, a committee can be used to “delegate” nodes to perform any kind of operation. In this work, to build and maintain our churn-resilient overlay network of committees we make use of the results in [13]. Our choice is motivated by two main advantages that the Spartan network in [13] has over the previous approach [29]: (a) it tolerates an abrupt adversarial churn rate of $\mathcal{O}(n/\log n)$, and (b) it can be built in $\mathcal{O}(\log n)$ rounds with high probability. Each committee is a dynamic random clique of size $\Theta(\log n)$ in which member nodes change continuously with the guarantee that the committee has $\Theta(\log n)$ nodes as its members at any given time. These committees are arranged into a wrapped butterfly network [53, 61]. The wrapped butterfly has $2^{k}$ rows and $k$ columns such that $k2^{k}\in\mathcal{O}(n/\log n)$ nodes and edges. The nodes correspond to committees and are represented by pairs $(r,\ell)$ where $\ell$ is the level or dimension of the committee ($0\leq\ell\leq k$) and $r$ is a $k$-bit binary number that denotes the row of the committee. Two committees $(r,\ell)$ and $(r^{\prime},\ell^{\prime})$ are linked by an edge that encodes a complete bipartite graph if and only if $\ell^{\prime}=\ell+1$ and either: (1) both committees are in the same row i.e., $r=r^{\prime}$, or (2) $r$ and $r^{\prime}$ differ in precisely the (+1(mod $(\ell+1(\text{mod }k))$th bit in their binary representation. Finally, the first and last levels of such network are merged into a single level. In particular, committee $(r,0)$ is merged into committee $(r,k)$. Figure 3 shows an example of Spartan network with $4$ rows and $2$ columns in which every node of the two-dimensional wrapped butterfly encodes a committee of $\Theta(\log n)$ random nodes and each edge between two committees encodes a complete bipartite graph connecting the vertices among committees. Notice that the first and last columns are the same set of committees since the butterfly is wrapped. The Spartan network is built during the bootstrap phase in which the adversary does not perform any move (having such phase is a common assumption in the DNC model [7, 8, 9, 10, 11, 13]). In [13], the authors showed how to efficiently build such wrapped butterfly of committees in $\mathcal{O}(\log n)$ rounds using $\mathcal{O}(\log n)$ messages per node at every round. For the sake of completeness, we provide a high-level description of how to build Spartan during the bootstrap phase. As a first step, a random node $v$ is elected as leader. Subsequently, a $\mathcal{O}(\log n)$ height binary tree rooted in $v$ is constructed. Next, the in-order traversal number for each node is computed and a cycle between the first $N=k2^{k}\in\mathcal{O}(n/\log n)$ nodes is created. Each node in such a cycle becomes a committee leader. The cycle is then transformed into a wrapped butterfly network with $k$ columns and $2^{k}$ rows. As a next step, the nodes that are not committee leaders randomly join one of the $N$ committees with the purpose to create committee of $\Theta(\log n)$ nodes. Once such committees are created, the nodes within each committee form a clique of size $\Theta(\log n)$. Finally, the committee leaders exchange the IDs of the nodes in their committee with their neighbors so that bipartite overlay edges can be formed between nodes in neighboring committees.
After the bootstrap phase, the nodes in Spartan run a continuous $\Theta(\log\log n)$ rounds maintenance cycle in which (1) all the nodes in the network move to another committee chosen uniformly at random among all the other committees and (2) the newly added nodes are assigned to random committees. This cycle prevents the powerful adversary to grasp sensitive information about the network structure and to replace all the nodes in the network. Augustine and Sivasubramaniam [13] showed that such a maintenance cycle makes the Spartan network robust with high probability against an adversarial churn rate of $\mathcal{O}(n/\log n)$, i.e., with high probability no committee is disrupted by the churn applied by oblivious adversary.
As a side contribution of our work, we generalize the Spartan protocol to handle an arbitrary number of nodes in the network, i.e., we relax the assumption that the size of the overlay network is fixed at each round to some number of nodes $n$. This makes the protocol adaptable to highly dynamic networks, in which the number of nodes might suddenly variate. We define a protocol that enlarges and reduces the dimensionality of $\mathcal{S}$ (i.e., reshapes $\mathcal{S}$) when the number of nodes in the Spartan network is such that the wrapped butterfly increases (resp. decreases) by one dimension. In Appendix C, we provide a reshaping protocol that reconfigures the Spartan network according to the number of peers present in the overlay network.
@@ 2.2 Replacing nodes that have been removed by the adversary
We describe how to maintain the live network $\mathcal{L}$ and the clean network $\mathcal{C}$ when churn occurs. Let us assume that a node $v\in V$ left the network. To preserve $\mathcal{C}$ and $\mathcal{L}$ structures, we require its committee in $\mathcal{S}$ “to cover” for the disappeared node. In other words, when a node $v$ leaves the network, its committee members in $\mathcal{S}$ will take care of all the operations involving $v$ in $\mathcal{L}$ and $\mathcal{C}$. Doing so, will temporarily preserve the structure of these networks, as if $v$ was still in the network. Moreover, we assume that at every round in the Spartan network, all the nodes within the same committee communicate their states (in $\mathcal{S},\mathcal{L}$, and $\mathcal{C}$) to all the other committee members, i.e., each node $u$ in a committee knows all its neighbors $N_{\mathcal{S}}(v)$, each neighbor in both skip list $\mathcal{L}$ and $\mathcal{C}$ for each $v\in N_{\mathcal{S}}(v)$, and all the committees of its neighbors in $\mathcal{L}$ and $\mathcal{C}$ [^7] The expected space needed to store the committee’s coordinates of $v$’ neighbors in $\mathcal{L}$ and $\mathcal{C}$ is $\mathcal{O}(\log n)$ because every node in a skip list has expected degree of $\mathcal{O}(\log n)$, w.h.p.. Furthermore, when a node $v$ leaves the network, each node in its committee in $\mathcal{S}$ will connect to $v$’s neighbors in $\mathcal{L}$ and $\mathcal{C}$. This can be done in constant time and will increase the degree of the nodes in the data structures to $\mathcal{O}(\text{polylog}(n))$. Notice that each committee in $\mathcal{S}$ is guaranteed to be robust against a $\mathcal{O}(n/\log n)$ churn rate, i.e., the probability that there exists some committee of size $o(\log n)$ is at least $1/n^{d}$ for an arbitrarily large $d\geq 1$ (see Theorem 6 in [12]). This implies that every committee of nodes in $\mathcal{S}$ is able to cover for the removed nodes in $\mathcal{L}$ and $\mathcal{C}$ with high probability. Figure 4 shows an example of how the committees in the Spartan network cover for its members that were removed by the adversary. More precisely, in Figure 4(b) we show how the orange node’s committee is covering for it after being dropped by the adversary. An alternative way to see this is to assume that the committee creates a temporary virtual red node in place of the orange one (see Figure 4(c)).
@@
Replacing node(s) that have been removed by the adversary requires $\mathcal{O}(1)$ number of rounds and work proportional to the churn.
@@
When a node $v$ leaves the network, all its committee neighbors create an edge with $v$ neighbors in $\mathcal{L}$ and $\mathcal{C}$. This requires $\mathcal{O}(1)$ rounds and the amount of edge that are created by $v$’s committee is $\mathcal{O}(\log^{2}n)$ (at most $4\log n$ edges for each node in the $\mathcal{O}(\log n)$-sized committee). Observe that in one cycle of $\mathcal{O}(\log n)$ rounds we can experience a churn of $\mathcal{O}(n/\log n)$ at every time instant. Thus, the overall work is $\mathcal{O}(n\cdot\text{polylog}{(n)})$ and respects our dynamic resource competitiveness constraints.
@@ 2.3 Deletion
We describe how to safely clean the network $\mathcal{C}$ in $\mathcal{O}(\log n)$ rounds from the removed nodes for which committees in the Spartan network $\mathcal{S}$ are temporarily covering for. The elements in the clean network $\mathcal{C}$ before the cleaning procedure can be of two types: (1) not removed yet and (2) patched-up, i.e., nodes that have been removed by the adversary and for which another group of nodes in the network is covering for. A trivial approach to efficiently remove the patched-up nodes from $\mathcal{C}$, would be the one of destroying the clean network $\mathcal{C}$ by removing these nodes and then reconstructing the cleaned skip list from scratch. However, this approach does not satisfy our dynamic resource competitiveness constraint (see Definition 1.1 and Section 1.3); the effort we must pay to build the skip list from scratch may be much higher than the adversary’s total cost. Hence, we show an approach to clean $\mathcal{C}$ that guarantees dynamic resource competitiveness. For the sake of explanation, assume that the patched-up nodes are red and the one that should remain in $\mathcal{C}$ are black. The protocol is executed in all the levels of the skip list in parallel. The high-level idea of the protocol is to create a tree rooted in the skip list’s left-topmost sentinel by backtracking from some “special” leaves located on each level of the skip list. Once the rooted tree is constructed, another backtracking phase in which we build new edges connecting black nodes separated by a contiguous for its leaves is executed. Figure 5 shows an example of the delete routine (Algorithm 2) run on all the levels of the skip list in Figure 5(a). Moreover, Figure 5(b) and Figure 5(c) show the execution of the delete routine at level 00. First, the tree rooted in the left-topmost sentinel is built by backtracking from the leaves (i.e., from the black nodes that have at least one red neighbor). Next, the tree is traversed again in a bottom-up fashion and green edges between black nodes separated by a list of red nodes are created (Figure 5(d)). Figure 5(e) depicts the skip list after after running Algorithm 2 in parallel on every level. Finally, Figure 5(f) shows the polished skip list.
@@
Figure 5: Example of the delete phase on a skip list. Red nodes are the one that must be deleted from the skip list, blue arrow indicates the path towards the left topmost sentinel in the skip list from the black nodes that wish wire-out its red neighbors. Figure 5(a) is the skip list before the deletion phase. Figure 5(b) shows the tree rooted in the left-topmost sentinel created by the Tree Formation Phase. Figure 5(c) shows the Message Initialization and the Propagation Phase executed at the bottom most level of the skip list where Blue circles indicate that an edge between two black nodes have been created. Figure 5(d) shows the skip list after the bottom most lever run Algorithm 2. Figure 5(e) is the skip list after running Algorithm 2 at all its levels. Finally, Figure 5(f) is the skip list after the deletion phase. The left and right sentinels are represented as ls and rs, respectively.
@@
To establish the forward direction of the bidirectional statement, let us consider the formation of an edge $(b,b^{\prime})$ at some node $u$ that is at a level $\ell^{\prime}\geq\ell$. It coalesced two pairs of the form $(x,\dot{b})$ and $(\dot{b^{\prime}},y)$. The implication is that $b$ has a right neighbor at level $\ell$ that is red and $\dot{b}$ has a left neighbor at level $\ell$ that is red. Moreover, they are the rightmost leaf and left most leaf respectively of the subtrees rooted at $u$’s children. Thus, none of the nodes between $b$ and $b^{\prime}$ at level $\ell$ are in $T$. This implies that they must all be red nodes.

### v2

@@ 2 Solution Architecture
Skip lists [67] are randomized data structures organized as a tower of increasingly sparse linked lists. Level 0 of a skip list is a classical linked list of all nodes in increasing order by key/ID. For each $i$ such that $i>0$, each node in level i1i-1 appears in level $i$ independently with some fixed probability $p$. The top lists act as “express lanes” that allow the sequence of nodes to be quickly traversed. Searching for a node with a particular key involves searching first in the highest level, and repeatedly dropping down a level whenever it becomes clear that the node is not in the current one. By backtracking on the search path it is possible to shows that no more than $\frac{1}{1-p}$ nodes are searched on average per level, giving an average search time of $\mathcal{O}(\log n)$ (see Appendix B for useful properties of randomized skip lists). We refer to the height of a skip list $\mathcal{L}$ (see Figure 3) as the maximum $h$ such that Level $h$ is not empty. Given a node $v$ in the skip list $\mathcal{L}$ with $n_{\mathcal{L}}$ elements, we use $N_{\mathcal{L}}^{(\texttt{right},\ell)}(v)$ and $N_{\mathcal{L}}^{(\texttt{left},\ell)}(v)$ to indicate $v$’s right and left neighbors at level $\ell$ respectively. Moreover, when the direction and the levels are not specified we refer to the overall set of neighbors of a node in a skip list, formally we refer to $N_{\mathcal{L}}(v)=\bigcup_{\ell\geq 0}(N_{\mathcal{L}}^{(\texttt{left},\ell)}(v)\cup N_{\mathcal{L}}^{(\texttt{right},\ell)}(v))$. Furthermore, we define $\ell_{\max}^{\mathcal{L}}(v)$ for a node $v\in\mathcal{L}$ to be its maximum height in the skip list [^1] We omit the superscript when the skip list is clear from the context. and $\ell_{\max}(\mathcal{L})=\max_{v\in\mathcal{L}}\ell_{\max}^{\mathcal{L}}(v)$ to be the overall height of the skip list.
Figure 3: Skip list with $n=6$ nodes and 33 levels.
@@
Moreover, in all skip lists, when a node exits the system, it is operated by a selected group of nodes. In the course of our algorithm description, if a node $u$ is required to perform some operation, but is no longer in the system, then its replacement node(s) will perform that operation on its behalf. Note further that some of the replacement nodes themselves may need to be replaced. Such replacement nodes will continue to represent $u$. The protocol assumes a short ($\Theta(\log n)$ round) initial “bootstrap” phase, where there is no churn [^2] Without a bootstrap phase, it is easy to show that the adversary can partition the network into large pieces, with no chance of forming even a connected graph. and it initializes the underlying network. More precisely, the bootstrap is divided in two sub-phases in which we (i) build the underlying churn resilient network described in Section 2.1 in $\mathcal{O}(\log n)$ rounds and, (ii) we build the skip lists data structures $\mathcal{L}$ and $\mathcal{C}$ (initially $\mathcal{L}=\mathcal{C}$) using the $\mathcal{O}(\log n)$ rounds technique described in Section 2.4; after this, the adversary is free to exercise its power to add or delete nodes up to the churn limit and the network will undergo a continuous maintenance process. The overall maintenance of the dynamic distributed data structure goes through cycles. Each cycle $c\geq 1$ in Algorithm 1 consists of four sequential phases. Without loss of generality, assume that initially $\mathcal{L}$ and $\mathcal{C}$ are the same. We use the notation $\mathcal{L}^{(c)}$ (resp. $\mathcal{S}^{(c)},\mathcal{C}^{(c)},\mathcal{B}^{(c)}$) to indicate the network $\mathcal{L}$ (resp. $\mathcal{S},\mathcal{C},\mathcal{B}$) during the cycle $c\in\mathbb{N}$, when it is clear from the context we omit the superscript to maintain a cleaner exposition.
@@
7 Phase 3 (Merge): The buffer $\mathcal{B}^{(c)}$ created in Phase 22 is merged into $\mathcal{C}^{(c)}$, i.e., $\mathcal{C}^{(c+1)}\leftarrow\mathcal{C}^{(c)}\cup\mathcal{B}^{(c)}$.
@@ 2.1 The Spartan network
A widely used approach to maintain a stable overlay network under high adversarial churn is to organize nodes into small-sized committees [7, 9, 29, 13]. A committee is a clique of small ($\Theta(\log n)$) size composed of essentially “random” nodes. A committee can be efficiently constructed, and more importantly, maintained under heavy churn. Maintenance must ensure that each committee is made up of a random set of $O(\log n)$ peers, thus making it difficult for an oblivious adversary to remove all nodes from a committee. Since each committee behaves as a single unit, the loss of a (strict) subset of peers, which an oblivious adversary can inflict, will not disrupt the committee; the remaining peers can ensure that the continues to operate as a single unit. Moreover, a committee can be used to “delegate” nodes to perform any kind of operation.
@@
In this work, we build on the Spartan network design [13], which improves on earlier committee-based approaches by (a) tolerating an adversarial churn rate of $\mathcal{O}(n/\log n)$ per round, and (b) achieving a construction time of $\mathcal{O}(\log n)$ rounds with high probability. Each committee is a dynamic random clique of size $\Theta(\log n)$ in which member nodes change continuously with the guarantee that the committee has $\Theta(\log n)$ nodes as its members at any given time. These committees are arranged into a wrapped butterfly network [53, 61]. The wrapped butterfly has $2^{k}$ rows and $k$ columns such that $k2^{k}\in\mathcal{O}(n/\log n)$ nodes and edges. The nodes correspond to committees and are represented by pairs $(r,\ell)$ where $\ell$ is the level or dimension of the committee ($0\leq\ell\leq k$) and $r$ is a $k$-bit binary number that denotes the row of the committee. Two committees $(r,\ell)$ and $(r^{\prime},\ell^{\prime})$ are linked by an edge that encodes a complete bipartite graph if and only if $\ell^{\prime}=\ell+1$ and either: (1) both committees are in the same row i.e., $r=r^{\prime}$, or (2) $r$ and $r^{\prime}$ differ in precisely the (+1(mod k))($\ell$+1($\text${mod }k))th bit in their binary representation. Finally, the first and last levels of such network correspond to the same level (i.e., the butterfly is wrapped). In particular, committees $(r,0)$, and $(r,k)$ are the same committees. An alternative way to represent this structure is to create edges from the k1k-1th level to the level 0333We decided to avoid this representation, to maintain a clear exposition and easy to read pictures.. Figure 4 shows an example of Spartan network with $4$ rows and $2$ columns in which every node of the two-dimensional wrapped butterfly encodes a committee of $\Theta(\log n)$ random nodes and each edge between two committees encodes a complete bipartite graph connecting the vertices among committees. Notice that the first and last columns are the same set of committees since the butterfly is wrapped.
The Spartan network is constructed during an initial bootstrap phase, assuming no adversarial activity [^4] Having such phase is a common assumption in the DNC model [7, 8, 9, 10, 11, 13], as follows: (1) a random leader node is elected; (2) a binary tree of height $\mathcal{O}(\log n)$ is built; (3) the first $N=k2^{k}$ nodes are organized into a cycle, and designated as committee leaders; (4) the cycle is transformed into a wrapped butterfly structure; (5) the remaining nodes are assigned randomly to committees, which then internally form cliques; (6) committee leaders exchange node ID lists to establish the bipartite overlay edges between committees.
After the bootstrap, Spartan maintains stability using a continuous maintenance cycle that runs every $\Theta(\log\log n)$ rounds. In each cycle, all nodes are reassigned uniformly at random to new committees, and any newly joined nodes are placed randomly as well. While each node can locally detect when a neighbor leaves (via the loss of the direct link), the system’s stability does not depend on explicitly tracking individual departures. Instead, the maintenance cycle ensures that, with high probability, each committee maintains $\Theta(\log n)$ members and the network remains connected, as long as the churn rate is $\mathcal{O}(n/\log n)$.
As a contribution of this work, we generalize the Spartan framework to handle variable network sizes, removing the assumption that the number of nodes remains fixed at $n$. We provide a reshaping protocol (Appendix C) that dynamically adjusts the dimensionality of the wrapped butterfly as the network size increases or decreases, allowing the system to maintain its structural and robustness guarantees under highly dynamic and more general conditions.
@@ 2.2 Replacing nodes that have been removed by the adversary
We describe how to maintain the live network $\mathcal{L}$ and the clean network $\mathcal{C}$ when churn occurs. Assume that node $v\in V$ left the network. To preserve $\mathcal{C}$ and $\mathcal{L}$ structures, we require its committee in $\mathcal{S}$ “to cover” for the disappeared node. In other words, when node $v$ leaves the network, its committee members in $\mathcal{S}$ will take care of all the operations involving $v$ in $\mathcal{L}$ and $\mathcal{C}$. This temporarily preserves the structure of these networks, as if $v$ were still present. Moreover, we assume that at every round in the Spartan network, all the nodes within the same committee communicate their states (in $\mathcal{S},\mathcal{L}$, and $\mathcal{C}$) to all the other committee members, i.e., each node $u$ in a committee knows all neighbors $N_{\mathcal{S}}(u)$ and all neighbors of these nodes in $\mathcal{L}$ and $\mathcal{C}$, and the committees they belong to [^5] The expected space needed to store the committee’s coordinates of $v$’ neighbors in $\mathcal{L}$ and $\mathcal{C}$ is $\mathcal{O}(\log n)$ because every node in a skip list has expected degree of $\mathcal{O}(\log n)$, w.h.p.. Furthermore, when a node $v$ leaves the network, each node in its committee in $\mathcal{S}$ will connect to $v$’s neighbors in $\mathcal{L}$ and $\mathcal{C}$. This can be done in constant time and will increase the degree of the nodes in the data structures to $\mathcal{O}(\text{polylog}(n))$. Notice that each committee in $\mathcal{S}$ is guaranteed to be robust against a $\mathcal{O}(n/\log n)$ churn rate, i.e., the probability that there exists some committee of size $o(\log n)$ is at most $1/n^{d}$ for an arbitrarily large $d\geq 1$ (see Theorem 6 in [12]). This implies that every committee of nodes in $\mathcal{S}$ is able to cover for the removed nodes in $\mathcal{L}$ and $\mathcal{C}$ with high probability. In particular, since each committee survives the churn with high probability (as guaranteed by Theorem 6  [12]), it cannot be fully destroyed. This in turn implies that the committee will succeed in covering the deleted node with high probability, ensuring the continuity of $\mathcal{L}$ and $\mathcal{C}$ despite adversarial removals. Figure 5 shows an example of how the committees in the Spartan network cover for its members that were removed by the adversary. More precisely, in Figure 5(b) we show how the orange node’s committee is covering for it after being dropped by the adversary. An alternative way to see this is to assume that the committee creates a temporary virtual red node in place of the orange one (see Figure 5(c)). This virtual node represents the collective responsibility of the committee, effectively simulating $v$’s presence.
@@
With high probability, replacing nodes removed by the adversary requires $\mathcal{O}(1)$ rounds and work proportional (within $\text{polylog}(n)$ factors) to the number of nodes removed.
@@
When a node $v$ leaves the network, its committee members in $\mathcal{S}$ take over its connections in $\mathcal{L}$ and $\mathcal{C}$ by creating direct edges to $v$’s neighbors in these structures. This takeover operation can be completed in $\mathcal{O}(1)$ rounds, as each committee member already knows the necessary local neighborhood information.
Each committee has expected size $\mathcal{O}(\log n)$, and each skip list node has expected degree $\mathcal{O}(\log n)$, so the total number of new edges created by $v$’s committee is at most $\mathcal{O}(\log^{2}n)$. Given that the system can experience up to $\mathcal{O}(n/\log n)$ churn events in one $\mathcal{O}(\log n)$-round cycle, the overall work is $\mathcal{O}(n\cdot\text{polylog}(n))$, satisfying dynamic resource competitiveness.
Crucially, by Theorem 6 in [12], each committee survives churn with high probability; that is, the probability that a committee’s size drops below $\mathcal{O}(\log n)$ is at most $1/n^{d}$ for arbitrarily large $d\geq 1$. Therefore, with high probability, the committee is able to successfully cover the departed node, ensuring the continuity of $\mathcal{L}$ and $\mathcal{C}$ despite adversarial removals.
@@ 2.3 Deletion
We describe how to safely clean the network $\mathcal{C}$ in $\mathcal{O}(\log n)$ rounds from the removed nodes for which committees in the Spartan network $\mathcal{S}$ are temporarily covering. The elements in the clean network $\mathcal{C}$ before the cleaning procedure can be of two types: (1) not removed yet and (2) patched-up, i.e., nodes that have been removed by the adversary and for which another group of nodes in the network is covering. A trivial approach to efficiently remove the patched-up nodes from $\mathcal{C}$, would be the one of destroying the clean network $\mathcal{C}$ by removing these nodes and then reconstructing the cleaned skip list from scratch. However, this approach does not satisfy our dynamic resource competitiveness constraint (see Definition 1.1 and Section 1.3); the effort we must pay to build the skip list from scratch may be much higher than the adversary’s total cost [^6] The adversary’s cost is the churn applied to the network, i.e., the number of nodes removed in a certain time period.. Hence, we show an approach to clean $\mathcal{C}$ that guarantees dynamic resource competitiveness. For the sake of explanation, assume that the patched-up nodes are red and the one that should remain in $\mathcal{C}$ are black. The protocol is executed in all the levels of the skip list in parallel. The high-level idea of the protocol is to create a tree rooted in the skip list’s left-topmost sentinel by backtracking from some “special” leaves located on each level of the skip list. Once the rooted tree is constructed, another backtracking phase in which we build new edges connecting black nodes separated by a chain of red nodes is executed. Figure 6 shows an example of the delete routine (Algorithm 2) run on all the levels of the skip list in Figure 6(a). Moreover, Figure 6(b) and Figure 6(c) show the execution of the delete routine at level 0. First, the tree rooted in the left-topmost sentinel is built by backtracking from the leaves (i.e., from the black nodes that have at least one red neighbor). Next, the tree is traversed again in a bottom-up fashion and green edges between black nodes separated by a list of red nodes are created (Figure 6(d)). Figure 6(e) depicts the skip list after running Algorithm 2 in parallel on every level. Finally, Figure 6(f) shows the skip list after the deletion phase.
@@
Figure 6: Example of the delete phase on a skip list. Red nodes are the one that must be deleted from the skip list, blue arrow indicates the path towards the left topmost sentinel in the skip list from the black nodes. Figure 6(a) is the skip list before the deletion phase. Figure 6(b) shows the tree rooted in the left-topmost sentinel created by the Tree Formation Phase. Figure 6(c) shows the Message Initialization and the Propagation Phase executed at the bottom most level of the skip list where blue circles indicate that an edge between two black nodes have been created. Figure 6(d) shows the skip list after the bottom most lever run Algorithm 2. Figure 6(e) is the skip list after running Algorithm 2 at all its levels. Finally, Figure 6(f) is the skip list after the deletion phase. The left and right sentinels are represented as ls and rs, respectively.
@@
To establish the forward direction of the bidirectional statement, let us consider the formation of an edge $(b,b^{\prime})$ at some node $u$ that is at a level $\ell^{\prime}\geq\ell$. It coalesced two pairs of the form $(x,\dot{b})$ and $(\dot{b^{\prime}},y)$. The implication is that $b$ has a right neighbor at level $\ell$ that is red and $b^{\prime}$ has a left neighbor at level $\ell$ that is red. Moreover, they are the rightmost leaf and left most leaf respectively of the subtrees rooted at $u$’s children. Thus, none of the nodes between $b$ and $b^{\prime}$ at level $\ell$ are in $T$. This implies that they must all be red nodes.